(a) For Vine 220 Vrms., determine the average value of the voltage delivered to the load by using a SCR with a firing angle of 100°. Also calculate the rms value of the voltage delivered to the load by using a Triac with a firing angle of 100° (b) What is the purpose of the resistor in thyristors snubbing circuit? What would be the effect if the resistor in the snubbing circuit is an open circuit?

Find: Δ.pp L, lpL, Avpp load, Po Pdíode P diode unre load switch driver pwm ーErelereac.
Answer #1

a)Vrms=220V

Vline=220 * 1.414 = 311 V in case of rectifier

Average output voltage  1 /2 TVin (1 + cos(a))

where firing angle->alpha

Therefo Vo= 49.46V

As we kn triac use in ac voltage regulators ,

Vo=Vm/pi sinalpha

=311/pi sin1000

=97.5 V

b)In the snubber circuit normally ployed for partial v/partial t protection

we use a capacitor in parrallel to prevent the turn on of the SCR due to high value of dv/dt as i= C dv/dt.Befor device is activated voltage builds across capacitor as it is charged and at the time on of device the voltage builds at sudden  across Sc.this may cause accidental turn on of scr and high current may damage device.R resistor in sereis with capacitor helps to conduct this current and consequently attenuate.